A body is thrown vertically upwards from the body of a cliff 100m high with a speed of 15m/s at the same time another body is dropped from the top of the cliff, when will the two bodies meet
1
Expert's answer
2020-09-23T09:00:15-0400

"v=v_0-gt\to t=v_0/g=15/9.81=1.53(s)"


"s_1=v_0t-gt^2\/2="


"=15\cdot1.53-9.81\cdot1.53^2/2=11.5(m)"


"s_2=gt^2/2=9.81\cdot1.53^2/2=11.5(m)"


So, under these conditions, the two bodies will meet on the surface of the Earth after


"t=\sqrt{2S/g}=\sqrt{2\cdot100/9.81}=4.51(s)"